Mr Sam spent 1/4 of his money on 7 pens and 9 highlighters. Each pen cost 3 times as much as a highlighter. He bought some more pens with 2/3 of his remaining money. How many pens did Mr Sam buy altogether?
Answer by greenestamps(13334)



He bought 7 pens and 9 highlighters; each pen cost 3 times as much as a highlighter. So what he spent was equivalent to 7(3)+9=30 highlighters, or 30/3=10 pens.

He spent 1/4 of his money on that purchase, so he has 3/4 of his money left. When he spent 2/3 of his remaining money on more pens, he spent 2/3 of 3/4 of his original amount of money.

2/3 of 3/4 of his original amount of money is 1/2 of his original money. That is twice as much as he spent on the first purchase (1/2 is twice as much as 1/4), so this second purchase was enough to buy 2*10=20 more pens.

So the total number of pens he bought was 7+20 = 27.

ANSWER: 27